How many ounces of pure salt must be added to 9oz of 5% salt solution to produce a 15% solution? (Pure salt is 100%.) Round your answer to the nearest hundredth of an oz}.
Answer by stanbon(75887)

How many ounces of pure salt must be added to 9oz of 5% salt solution to produce a 15% solution?
-----
Equation::
salt + salt = salt
1*x + 0.05*9 = 0.15(x+9)
100x + 5*9 = 15x + 15*9
85x = 10*9
x = 1.06 ounces (amt. of pure salt to add)
----------------
Cheers,
Stan H.
------------------